Hi. I did these two problems and I just wanted to make sure I did them correctly.
Question:
Write the slope intercept form of the equation of the line passing through the two points (2,-3) and (-1,-1)
My answer is...y=-2/3x-5/3
2nd question:
write the slope intercept form of the equation of the line satisfying the given conditions: parallel to 3x+4y=-1 and through (3,-2)
My answer is ....y=-3/4x+1/4

are my answers correct? If not, what did I do wrong? Thanks so much!

Answer by Alan3354(69443)

Hi. I did these two problems and I just wanted to make sure I did them correctly.
Question:
Write the slope intercept form of the equation of the line passing through the two points (2,-3) and (-1,-1)
My answer is...y=-2/3x-5/3
2nd question:
write the slope intercept form of the equation of the line satisfying the given conditions: parallel to 3x+4y=-1 and through (3,-2)
My answer is ....y=-3/4x+1/4

are my answers correct? If not, what did I do wrong? Thanks so much!
==================
You can check them. Sub the x & y values of the points, see if they fit.
---
(2,-3) and (-1,-1)
My answer is...y=-2/3x-5/3
----
-3 =? -(2/3)*2 - 5/3
-3 =? -4/3 - 5/3 = -9/3 = -3
That point is on the line.
---
-1 =? (-2/3)*-1 - 5/3
-1 =? 2/3 - 5/3 = -1
Both points fit the equation, so it's correct.
---------
Check the other one the same way.
The slope is correct, check the point.